The difference of the effective atomic number and the number of unpaired electrons of Fe in
K4[Fe(CN6)]=x.Findthevalueofx/4

Grade:12

1 Answers

Vikas TU
4 years ago
Dear student 
EAN means metal atomic no. Minus oxidation no.of metal atoms plus 2× coordination number,which is equal to next atomic number.
In k4(fe(cn)6) , central atom Fe. I..e, 26.
Oxidation no.=4(x(-1)6) =2
So Fe is +2
Coordination no=2×6 =12
Then EAN =26–2×12 =36
 the 4 unpaired d-electrons , making the co-ordination compound diamagnetic hence , has 0 unpaired electrons
So x = 36 
x/4 =9